OpenStudy (anonymous):

Helpp! I can't think of anything!! Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 10 3x + 4y = 25

OpenStudy (ganpat):

Part 1: say x is number of Oranges and y is number of Apples .. You & ur friend go to market to buy them you buy 2 orange and a apple at a rate of 10.. and your friend buys 3 oranges and 4 apples for a rate of 25..

OpenStudy (ganpat):

Part 2: Now you have to find the exact amount for each orange and apple.. So solve them simultaneously.. Multiply eqnt 1 by 4 8x + 4y = 40 3x + 4y = 25 5x = 15... (subtracting the equations.. ) x = 3... price for single Orange.. substitute in any of the equation.. 2 (3) + y = 10 y = 4.. price of single apple So price for one orange = 3 and for a apple = 4...

OpenStudy (anonymous):

Thank you! That helped a lot!

OpenStudy (ganpat):

w/c... Just write in ur own words.. speak whatever u write..

OpenStudy (anonymous):

Okay